Skip to content

OSG Technology Area Meeting, 23 July 2018

Coordinates: Conference: 719-284-5267, PIN: 57363; https://www.uberconference.com/osgblin
Attending: BrianL, Carl, Derek, Edgar, Mat, Suchandra

Announcements

  • TimT OOO, returning Wednesday
  • BrianL out next Monday

Triage Duty

  • This week: Edgar
  • Next week: Mat
  • 7 (+0) open tickets

JIRA

# of tickets Δ State
135 +3 Open
16 -3 In Progress
5 +3 Ready for Testing
0 +0 Ready for Release

OSG Software Team

  • OSG 3.4.16: XRootD 4.8.4, Frontier Squid 3.5.27-5.1, and SciTokens
  • TimC offered his assistance with any web development
  • Review the next triage rotation
  • Let's get the GCT tarballs onto the repo host this week
  • Next doc focus 8/2 1PM CDT

Discussion

  • BrianL (AI) - promote SciTokens ticket to testing, assign squid ticket to Suchandra
  • Derek (AI) - Make changes to repo.gridcf.org for the GCT tarballs

Support Update

  • LIGO (Edgar) - Some sites don't have secure CVMFS installed (it is difficult for site admins to test)
  • OSU (BrianL, Mat) - inaccurate site topology and still running a GRAM gatekeeper
  • GridUNESP (BrianL) - investigating N/A site records in the GRACC reporting that's suspected to be GridUNESP sites

OSG Release Team

3.4.16 Δ Status
4 -5 Open
0 -1 In Progress
4 +3 Ready for Testing
0 +0 Ready for Release
8 -3 Total
  • OSG 3.4.16
    • Testing
      • XRootD 4.8.4
      • Frontier Squid
    • Ready for Release
      • Nothing Yet
  • Data
    • Testing
      • cilogon-openid-ca
      • vo-client-dcache
    • Ready for Release
      • Nothing yet
  • GOC
    • Testing
      • Drop mirror.batlab.org from mirror list
    • Ready for Release
      • Nothing yet

Discussion

None this week

OSG Investigations Team

Operations

  • Moving hosts from Anvil cloud host at Nebraska for Anvil downtime
    • Lots of hosts moved already. Hopefully, without any noticable downtime
    • GRACC head node is moving today (data nodes were moved 2 weeks ago). May see small downtimes (but will be up at the end of the day)
    • Was able keep the same IPs for the central collectors, CE's automatically picked up the (non)change and reported to the new collectors.
  • Need to start operations meetings again.

TODO:

  • Discussion of SLA's, availability metrics
  • Puppetize: ~~Collectors~~, GRACC

Investigations

In Progress:

  • Discussions with with Internet2 to place StashCache in connection points.

Ongoing

Discussions

None this week